# Dwarf Galaxies in the MATLAS Survey: Hubble Space Telescope Observations of Nuclear Star Clusters

### v1

@@ 6.3 NSCs-to-galaxy stellar mass relation
Several studies have reported correlations between some properties of the NSCs and the host galaxies, especially their luminosity and mass, in both dwarf and more massive galaxies (e.g., Balcells et al. 2003; Graham & Guzmán 2003; Böker et al. 2004; Turner et al. 2012; Scott & Graham 2013; Poulain et al. 2021; Carlsten et al. 2022; Hoyer et al. 2023). In this section we focus on the NSC-to-galaxy mass ratio. This relation has been studied for a large range of galaxies and environment, and thus is well suited for a comparison with our sample. We show at the top of Figure 4 the relation between the stellar mass of the NSC and that of its host galaxy, while we plot the NSC-to-galaxy mass ratio as a function of the galaxy stellar mass in the bottom part. We compare the HST NSCs to those observed in the LV, in the Fornax and Virgo clusters, as well as in LTGs, using all the samples with available stellar masses for both the NSCs and hosts described in Section 6.1.
@@
In agreement with the literature, we observe an overall increase of the mass of the NSC with that of the host. We study this increase by computing the running average of the NSC mass per bin of galaxy stellar mass, combining all the samples. Interestingly, we find that a majority of the NSCs in the HST sample falls above the running average, suggesting that the NSCs tend to be more massive in our sample. While these NSCs appear to also deviate from the relation fitted by Neumayer et al. (2020) on dwarf galaxies, considering the error bars on the MATLAS dwarfs, they mostly agree within the estimate uncertainties on the relation fitted on dEs from the Coma cluster (Graham & Guzmán 2003). We investigate the deviant galaxies by looking in details at the properties of the dwarf galaxies with high and mass ratio in the Virgo cluster. We find that they tend to either have a very high nuclear-to-total luminosity fraction or are UDGs. According to Wang et al. (2023), all have properties of UCD progenitors and are ongoing tidal disruption. The possible evolution of the MATLAS nucleated galaxies observed with HST into UCDs is discussed in Section 8. Another possible explanation could be that some of these galaxies might have experienced more GC mergers than other dwarfs due to a larger population of GCs, leading to more massive NSCs (Turner et al. 2012). Over the past decade, UDGs with a large number of GCs were identified (e.g., Lim et al. 2020; Müller et al. 2021; Forbes & Gannon 2024; Marleau et al. 2024).
@@
We note that we observe an overall change in the relation at a galaxy stellar mass of about in the lower panel of Figure 4, already mentioned in previous works (Ordenes-Briceño et al. 2018; Sánchez-Janssen et al. 2019; Neumayer et al. 2020; Su et al. 2022). The running average of low-mass galaxies follows the fitted relation of Neumayer et al. (2020), while the massive galaxies have an almost flat relation, similar to the one fitted by Neumayer et al. (2020) using the massive galaxies with dynamical and spectroscopic measurements for the NSC mass from Erwin & Gadotti (2012), and the relation fitted for ETGs and dEs in Scott & Graham (2013). Part of this change in the relation could be due to a change of dominant morphology of the galaxies: ETG for the low-mass sample, and LTG for the high-mass sample. The effect of the morphology on the NSCs-to-galaxy stellar mass relation has been investigated in several studies (Georgiev et al. 2016; Neumayer et al. 2020). In addition, Scott & Graham (2013) reported a steepening of the relation when the sample contains nuclear disk. While the presence of such disk was not reported in the samples used here, it might explain the deviation from the relation of some galaxies with .
@@ 6.4 Color of the NSCs
In this section, we study the stellar populations of the NSCs by looking at their integrated color as well as the variation of the color with radius, i.e., their color profile. The modeled NSCs show a range of color indices from 0.54 to 1.23 with a mean mag, similar to the observed colors of GCs in the MATLAS UDGs (Müller et al. 2021; Marleau et al. 2024). No trends are observed between the color of the NSCs and its absolute magnitude, the galaxy color, or the galaxy stellar mass.
We defined the color profiles of the NSCs as the difference between the light profiles in the F606W and F814W images that we produced making use of autoprof (Stone et al. 2021) fixing the same center and ellipses in both bands. We selected only the 15 marginally resolved NSCs with Re,NSC¿1 pix to ensure having multiple pixels. We display the color profiles within 3Re,NSC ordered by increasing M in Figure 5, and the individual profiles in Figure 10, to investigate a possible change of the profile with the NSC mass. We observe that a large majority of the low-mass NSCs, with M M⊙, tend to have an almost constant profile, while the massive ones show a change of color, with a red outer area and an inner region that gets bluer towards the photocenter. We note that both types of profile can be found at M from to M⊙.
The absence of color gradient in the center of the low-mass NSCs means there is an homogeneous stellar population, arguing in favor of GC merging. On the contrary, the change of color can be explained by a change of stellar population. In the high-mass-end of NSCs in dwarf galaxies, a combination of formation scenario has been reported (Fahrion et al. 2021, 2022a, 2022b). Such composite scenario matches our observations, starting with GC infall producing the outer region red color, and followed by in-situ star formation causing a younger, bluer, stellar population at the center. Given the low spatial resolution of the NSCs, the observed trend needs to be confirmed by observing more NSCs, especially in nearby galaxies.
An exception to this color trend is the low-mass NSC of MATLAS-49 with a of 5.91. The latter has an opposite profile to the massive NSCs, with a blue stellar population in the outer region that gets redder towards the photocenter. In the context of dwarf galaxies, this type of profile could be due to circumnuclear star formation, as observed for some NSCs in dEs (Paudel & Yoon 2020) following the wet migration scenario predictions. According to this scenario, MATLAS-49 should be a gas-rich dE, but no HI-detection was reported for this galaxy in Poulain et al. (2022). Given its location we estimate its HI mass to be less than M⊙.
@@ 7 Formation of NSCs from GCs
The GC population of the MATLAS UDGs observed by HST has been studied in detail in Marleau et al. (2024). Making use of this work, we first investigate the radial distribution of the GCs as a function of their stellar mass to question the effect of dynamical friction. In a second part, we look at the Re – luminosity relation of both the GCs and NSCs in the context of the migration scenario.
@@ 7.1 GC radial distribution
The strength of the dynamical friction depends on the steepness of the density profile of the galaxy, but also on the mass of the GC and its initial position (Hernandez & Gilmore 1998; Lotz et al. 2001; Arca-Sedda & Capuzzo-Dolcetta 2014; Arca-Sedda et al. 2015), so that the most massive GCs, and GCs at small radial distances, are the fastest to migrate to the center and the most likely to merge to form a NSC, and galaxies with the steepest profiles have the shortest orbital decay timescales. To explore this GC mass segregation, we look at the radial distance of the GCs of the nucleated and non-nucleated dwarfs as a function of their mass. We only modeled the marginally resolved GC candidates detected in the 17 dwarfs located below 25 Mpc, as the GC systems are dominated by unresolved sources at greater distances (see Marleau et al. 2024 for further details). Following the same approach to the NSCs described in Section 5, we obtained a reliable Sérsic model for 142 GCs in at least one of the two filters, from which we derived the colors and stellar masses. The properties of the GCs are available in Table 2, and we indicate the GCs with R pix, as for the NSCs.
@@
In Figure 6, we show the dependence of the radial distribution on the stellar mass of the GCs within two 2Re,gal of the nucleated and non-nucleated dwarfs, together with the running average of the radial distance per bin, 0.5 log(M/M⊙). We looked for an effect of the Sérsic index of the host galaxy on the radial distribution of the GCs using the structural properties from Poulain et al. (2021). We find that the nucleated galaxies have, on average, a higher Sérsic index than the non-nucleated dwarfs with 1.01 and 0.85, respectively, suggesting a steeper density profile for the nucleated dwarfs and a more efficient dynamical friction. However, the galaxies show a wide range of Sérsic index, from about 0.5 to 1.6, regardless of the presence of a NSC, and we do not find any influence of the Sérsic index on the radial distribution of the GCs. We observe that GCs in both nucleated and non-nucleated galaxies have a similar radial distribution for M¡ 10M⊙, with the GCs being located at all –as shown by the large standard deviation on the running average– except for GCs with M¿ 10M⊙, which are not found beyond 1.5 Re,gal in the non-nucleated galaxies. We note that the sample is rather incomplete on the low-mass-end, since we consider here only marginally resolved GCs. Thus, the radial distribution of GCs with M¡ 10M⊙ cannot be assessed. Considering the most massive GCs with M¿ 10M⊙, we find opposite trends: the GCs are mostly found within 0.5 Re,gal in non-nucleated dwarfs, while they are located above 1 Re,gal in nucleated dwarfs. These tendencies are based on low statistics, but might highlight the effect of dynamical friction, where the most massive GCs have migrated to the center of the non-nucleated dwarf, while the most massive GCs in the nucleated might have merged to form the NSCs, hence the lack of these GCs in the central region similar to the findings of Lotz et al. (2001) for nucleated dEs in the Virgo cluster.
@@ 7.2 Re – luminosity relation
The simulations of NSC formation from GC mergers undertaken in Bekki et al. (2004) have shown that the relationship between the structural properties and the luminosity of the NSCs differ from those found for GCs. We investigated in Section 6.2 the relation between Re and MV of the MATLAS NSCs. We now compare this scaling relation to that of the MATLAS GCs. In Figure 7, we plot Re as a function of MV for both types of star clusters, and compare the MATLAS GCs to those in the Milky Way (MW; Harris 1996, 2010 edition) and M31 (Peacock et al. 2010). Unsurprisingly, the MATLAS GCs occupy a similar range of MV to the MW and M31 GCs. In agreement with Bekki et al. (2004), we observe a different trend for the GCs than for the NSCs, where the maximum Re reached by the GCs at a given MV decreases towards brighter GCs, while it is constant for the NSCs. Hoyer et al. (2023) noted that the NSC outliers in the LV seem to follow the size-mass and size-luminosity relations observed for the GCs. More statistics for low-mass NSCs is needed to assess if they they fall into the GC regime.
@@ 8 Nucleated dwarfs as UCD progenitors
UCDs are a class of galaxies as bright as dwarf galaxies but more compact, in between the size of dwarfs and GCs. One of their formation scenarios implies that UCDs formed from the stripped nuclei remnants of tidally disrupted nucleated dwarf galaxies (e.g., Zinnecker et al. 1988; Freeman 1993; Bekki et al. 2001). Simulations argue that this is the predominant formation path of the most massive UCDs (Pfeffer et al. 2014; Mayes et al. 2021). Moreover, this scenario is also supported by the observation of tidal features around some UCDs (Voggel et al. 2016; Schweizer et al. 2018), asymmetric or elongated shapes (Wittmann et al. 2016), diffuse haloes (Evstigneeva et al. 2008; Liu et al. 2020), similarities between the stellar populations of UCDs and NSCs (Paudel et al. 2010; Norris et al. 2015; Janz et al. 2016; Wang et al. 2023), and the presence of MBHs in UCDs (Graham 2020; Mayes et al. 2024). An example of UCD candidate in the process of formation has already been identified in the MATLAS dwarf galaxy sample. MATLAS-167 is located at the end of a tidal tail extending from the giant S0 galaxy NGC936; a detailed analysis of the object can be found in Paudel et al. (2023).
@@
We investigate whether some of the NSCs in our HST sample could be UCD progenitors. In Figure 7, we compare the Re – luminosity scaling relation of the modeled MATLAS NSCs to the UCDs from the Virgo and Fornax galaxy clusters (Liu et al. 2020; Saifollahi et al. 2021). We split the Virgo sample to highlight the UCDs with identified outer envelopes, and we only show UCDs with R pc in the Fornax sample, as Saifollahi et al. (2021) question the reliability of the measurement of smaller sizes. During the formation of UCDs, up to 98% of the stellar mass of the MATLAS dwarfs will get stripped (Bekki et al. 2001; Wang et al. 2023), leaving the NSCs slightly larger and brighter than they currently are. As the MATLAS NSCs appear to form a bridge between the GCs and UCDs, we suggest that they could be UCD progenitors if their dwarf galaxy host undergoes tidal disruption.
@@
The study of Wang et al. (2023) presents different phases of UCD formation from tidally stripped dE,N in the Virgo cluster. They suggest that, during the first transitional phase, tidal stripping can remove more than 90% of the galaxy stellar mass, creating so-called strongly nucleated dwarfs with a nuclear-to-total-luminosity fraction reaching . In Figure 8, we show the NSC-to-galaxy luminosity relation of our HST sample, with the nuclear-to-total luminosity fraction of the NSCs as a colorbar. The NSCs show overall a good agreement with the relation derived by Graham & Guzmán (2003). We investigate if the five dwarfs outliers with a fraction show signs of tidal disruption. We identify tidal features from nearby massive galaxies related to two of them, MATLAS-1539 and MATLAS-1577, which have a fraction of 6.6% and 9.2%, respectively. The dwarfs and corresponding tidal features are visible in Figure 9. MATLAS-1539 seems compact and is located on top of a tidal tail from NGC5198, while MATLAS-1577 has a disturbed morphology and is extended towards a stellar stream from NGC5308. We note that the NSC of MATLAS-1539 is the largest of the modeled ones, with a Re,NSC of 21.58 pc, and already follows a similar MV–Re relation to the UCDs (see Figure 7). Following these results, even though only MATLAS-1577 meets the criteria of strongly nucleated dwarfs from Wang et al. (2023), we suggest that both MATLAS-1539 and MATLAS-1577 are candidate UCD progenitors. Among the galaxies with a nuclear-to-total luminosity fraction below 6%, we found three tidally elongated UDGs located close to massive galaxies. According to Wang et al. (2023), at early steps of tidal disruption, nucleated dEs fall into the regime of UDGs and show tidal features such as and extended stellar envelope or tidal tails. Based on this statement, we add these three UDGs to the list of candidates, and show them in Figure 9.

### v2

@@ 6.3 NSC-to-galaxy stellar mass relation
Several studies have reported correlations between some properties of the NSCs and the host galaxies, especially their luminosity and mass, in dwarf galaxies and in more massive galaxies (e.g., Balcells2003; Graham2003; Boeker2004; Turner2012; Scott2013; Poulain2021; Carlsten2022; Hoyer2023). In this section we focus on the NSC-to-galaxy mass ratio. This relation has been studied for a large range of galaxies and environments, and thus is well suited for a comparison with our sample. We show at the top of Fig. 4 the relation between the stellar mass of the NSC and that of its host galaxy, while we plot the NSC-to-galaxy mass ratio as a function of the galaxy stellar mass in the bottom part. We compare the HST NSCs to those observed in the LV, Fornax, and Virgo clusters, and in LTGs, using all the samples with available stellar masses for both the NSCs and hosts described in Sect. 6.1.
@@
In agreement with the literature, we observe an overall increase in the mass of the NSC with that of the host. We studied this increase by computing the running average of the NSC mass per bin of galaxy stellar mass, combining all the samples. Interestingly, we find that a majority of the NSCs in the HST sample fall above the running average, suggesting that the NSCs tend to be more massive in our sample. While these NSCs appear to also deviate from the relation fitted by Neumayer2020 on dwarf galaxies, considering the error bars on the MATLAS dwarfs, they mostly agree within the estimate uncertainties on the relation fitted on dEs from the Coma cluster (Graham2003). We investigated the deviant galaxies by looking in detail at the properties of the dwarf galaxies with high and mass ratio in the Virgo cluster. We find that either they tend to have a very high nuclear-to-total luminosity fraction or they are UDGs. According to Wang2023, all have properties of UCD progenitors and are undergoing tidal disruption. The possible evolution of the MATLAS nucleated galaxies observed with HST into UCDs is discussed in Sect. 8. Another possible explanation could be that some of these galaxies might have experienced more GC mergers than other dwarfs due to a larger population of GCs, leading to more massive NSCs (Turner2012). Over the past decade, UDGs with a large number of GCs have been identified (e.g., Lim2020; Mueller2021; forbes2024; Marleau2024).
@@
We note that we observe an overall change in the relation at a galaxy stellar mass of about in the lower panel of Fig. 4, already mentioned in previous works (Ordenes2018; Janssen2019; Neumayer2020; Su2022). The running average of low-mass galaxies follows the fitted relation of Neumayer2020, while the massive galaxies have an almost flat relation, similar to the one fitted by Neumayer2020 using the massive galaxies with dynamical and spectroscopic measurements for the NSC mass from Erwin2012, and the relation fitted for ETGs and dEs in Scott2013. Part of this change in the relation could be due to a change in the dominant morphology of the galaxies: ETG for the low-mass sample and LTG for the high-mass sample. The effect of the morphology on the NSC-to-galaxy stellar mass relation has been investigated in several studies (Georgiev2016; Neumayer2020). In addition, Scott2013 reported a steepening of the relation when the sample contains a nuclear disk. While the presence of such a disk was not reported in the samples used here, it might explain the deviation from the relation of some galaxies with .
@@ 6.4 Color of the NSCs
In this section we study the stellar populations of the NSCs by looking at their integrated color as well as the variation of the color with radius, i.e., their color profile. The modeled NSCs show a range of color indices from 0.54 to 1.23 with a mean mag, similar to the observed colors of GCs in the MATLAS UDGs (Mueller2021; Marleau2024). No trends are observed between the color of the NSCs and its absolute magnitude, the galaxy color, or the galaxy stellar mass.
We defined the color profiles of the NSCs as the difference between the light profiles in the F606W and F814W images that we produced making use of autoprof (Stone2021b) fixing the same center and ellipses in both bands. We selected only the 15 marginally resolved NSCs with Re,NSC¿1 pix to ensure there were multiple pixels. We display the color profiles within 3Re,NSC ordered by increasing M in Fig. 5, and the individual profiles in Fig. 10, to investigate a possible change in the profile with the NSC mass. We observe that a large majority of the low-mass NSCs, with M M⊙, tend to have an almost constant profile, while the massive ones show a change in color: a red outer area and an inner region that gets bluer toward the photocenter. We note that both types of profile can be found at M from to M⊙.
The absence of a color gradient in the center of the low-mass NSCs means there is a homogeneous stellar population, arguing in favor of GC merging. On the contrary, the change in color can be explained by a change in stellar population. At the high-mass end of NSCs in dwarf galaxies, a combination of formation scenarios has been reported (Fahrion2021; Fahrion2022b; Fahrion2022). This composite scenario matches our observations, starting with GC infall producing the outer region red color, followed by in situ star formation causing a younger bluer stellar population at the center. Given the low spatial resolution of the NSCs, the observed trend needs to be confirmed by observing more NSCs, especially in nearby galaxies.
An exception to this color trend is the low-mass NSC of MATLAS-49 with a of 5.91. This NSC has an opposite profile to the massive NSCs, with a blue stellar population in the outer region that gets redder toward the photocenter. In the context of dwarf galaxies, this type of profile could be due to circumnuclear star formation, as observed for some NSCs in dEs (Paudel2020) following the wet migration scenario predictions. According to this scenario, MATLAS-49 should be a gas-rich dE, but no HI-detection was reported for this galaxy in Poulain2022. Given its location, we estimate its HI mass to be less than M⊙.
@@ 7 Formation of NSCs from GCs
The GC population of the MATLAS UDGs observed by HST is studied in detail in Marleau2024. Making use of this work, we first investigated the radial distribution of the GCs as a function of their stellar mass to question the effect of dynamical friction. In a second part, we look at the Re – luminosity relation of the GCs and the NSCs in the context of the migration scenario.
@@ 7.1 GC radial distribution
The strength of the dynamical friction depends on the steepness of the density profile of the galaxy, but also on the mass of the GC and its initial position (Hernandez1998; lotz2001; Arca-Sedda2014; Arca-Sedda2015), so that the most massive GCs, and GCs at small radial distances, are the fastest to migrate to the center and are the most likely to merge to form a NSC, and galaxies with the steepest profiles have the shortest orbital decay timescales. To explore this GC mass segregation, we look at the radial distance of the GCs of the nucleated and non-nucleated dwarfs as a function of their mass. We only modeled the marginally resolved GC candidates detected in the 17 dwarfs located below 25 Mpc, as the GC systems are dominated by unresolved sources at greater distances (see Marleau2024 for further details). Following the same approach to the NSCs described in Sect. 5, we obtained a reliable Sérsic model for 142 GCs in at least one of the two filters, from which we derived the colors and stellar masses. The properties of the GCs are available in Table 2, and we indicate the GCs with R pix, as for the NSCs.
@@
In Fig. 6 we show the dependence of the radial distribution on the stellar mass of the GCs within 2Re,gal of the nucleated and non-nucleated dwarfs, together with the running average of the radial distance per bin, 0.5 log(M/M⊙). We looked for an effect of the Sérsic index of the host galaxy on the radial distribution of the GCs using the structural properties from Poulain2021. We find that the nucleated galaxies have, on average, a higher Sérsic index than the non-nucleated dwarfs with 1.01 and 0.85, respectively, suggesting a steeper density profile for the nucleated dwarfs and a more efficient dynamical friction. However, the galaxies show a wide range of Sérsic index, from about 0.5 to 1.6, regardless of the presence of a NSC, and we do not find any influence of the Sérsic index on the radial distribution of the GCs. We observe that GCs in both nucleated and non-nucleated galaxies have a similar radial distribution for M¡ 10M⊙, with the GCs being located at all –as shown by the large standard deviation on the running average– except for GCs with M¿ 10M⊙, which are not found beyond 1.5 Re,gal in the non-nucleated galaxies. We note that the sample is incomplete at the low-mass end since we consider here only marginally resolved GCs. Thus, the radial distribution of GCs with M¡ 10M⊙ cannot be assessed. Considering the most massive GCs with M¿ 10M⊙, we find opposite trends: the GCs are mostly found within 0.5 Re,gal in non-nucleated dwarfs, while they are located above 1 Re,gal in nucleated dwarfs. These tendencies are based on low statistics, but might highlight the effect of dynamical friction, where the most massive GCs have migrated to the center of the non-nucleated dwarf, while the most massive GCs in the nucleated might have merged to form the NSCs. Hence, the lack of these GCs in the central region is similar to the findings of lotz2001 for nucleated dEs in the Virgo cluster.
@@ 7.2 Re – luminosity relation
The simulations of NSC formation from GC mergers undertaken in Bekki2004 have shown that the relationship between the structural properties and the luminosity of the NSCs differ from those found for GCs. We investigated in Sect. 6.2 the relation between Re and MV of the MATLAS NSCs. We now compare this scaling relation to that of the MATLAS GCs. In Fig. 7 we plot Re as a function of MV for both types of star clusters, and compare the MATLAS GCs to those in the MW citepHarris1996, 2010 edition) and M31 (Peacock2010). Unsurprisingly, the MATLAS GCs occupy a similar range of MV to the MW and M31 GCs. In agreement with Bekki2004, we observe a different trend for the GCs than for the NSCs, where the maximum Re reached by the GCs at a given MV decreases toward brighter GCs, while it is constant for the NSCs. Hoyer2023 noted that the NSC outliers in the LV seem to follow the size-mass and size-luminosity relations observed for the GCs. More statistics for low-mass NSCs is needed to assess whether they fall into the GC regime.
@@ 8 Nucleated dwarfs as UCD progenitors
Ultra-compact dwarf galaxies are a class of galaxies as bright as dwarf galaxies but more compact, in between the size of dwarfs and GCs. One of their formation scenarios implies that UCDs formed from the stripped nuclei remnants of tidally disrupted nucleated dwarf galaxies (e.g., Zinnecker1988; Freeman1993; Bekki2001). Simulations argue that this is the predominant formation path of the most massive UCDs (Pfeffer2014; Mayes2021). Moreover, this scenario is also supported by the observation of tidal features around some UCDs (Voggel2016; Schweizer2018), asymmetric or elongated shapes (Wittmann2016), diffuse halos (Evstigneeva2008; Liu2020), similarities between the stellar populations of UCDs and NSCs (Paudel2010; Norris2015; Janz2016; Wang2023), and the presence of MBHs in UCDs (Graham2020; Mayes2024). An example of a UCD candidate in the process of formation has already been identified in the MATLAS dwarf galaxy sample. MATLAS-167 is located at the end of a tidal tail extending from the giant S0 galaxy NGC936; a detailed analysis of the object can be found in Paudel2023.
@@
We investigated whether some of the NSCs in our HST sample could be UCD progenitors. In Fig. 7 we compare the Re – luminosity scaling relation of the modeled MATLAS NSCs to the UCDs from the Virgo and Fornax galaxy clusters (Liu2020; Saifollahi2021). We split the Virgo sample to highlight the UCDs with identified outer envelopes, and we only show UCDs with R pc in the Fornax sample, as Saifollahi2021 question the reliability of the measurement of smaller sizes. During the formation of UCDs, up to 98% of the stellar mass of the MATLAS dwarfs will get stripped (Bekki2001; Wang2023), leaving the NSCs slightly larger and brighter than they currently are. As the MATLAS NSCs appear to form a bridge between the GCs and UCDs, we suggest that they could be UCD progenitors if their dwarf galaxy host undergoes tidal disruption.
@@
The study of Wang2023 presents different phases of UCD formation from tidally stripped nucleated dEs in the Virgo cluster. They suggest that, during the first transitional phase, tidal stripping can remove more than 90% of the galaxy stellar mass, creating strongly nucleated dwarfs with a nuclear-to-total-luminosity fraction reaching . In Fig. 8 we show the NSC-to-galaxy luminosity relation of our HST sample; the nuclear-to-total luminosity fraction of the NSCs is shown as a color bar. The NSCs show an overall good agreement with the relation derived by Graham2003. We investigated whether the five dwarfs outliers with a fraction show signs of tidal disruption. We identified tidal features from nearby massive galaxies related to two of them, MATLAS-1539 and MATLAS-1577, which have a fraction of 6.6% and 9.2%, respectively. The dwarfs and corresponding tidal features are visible in Fig. 9. MATLAS-1539 seems compact and is located on top of a tidal tail from NGC5198, while MATLAS-1577 has a disturbed morphology and is extended toward a stellar stream from NGC5308. We note that the NSC of MATLAS-1539 is the largest of the modeled ones, with a Re,NSC of 21.58 pc, and already follows a similar MV–Re relation to the UCDs (see Fig. 7). Following these results, even though only MATLAS-1577 meets the criteria of strongly nucleated dwarfs from Wang2023, we suggest that MATLAS-1539 and MATLAS-1577 are both candidate UCD progenitors. Among the galaxies with a nuclear-to-total luminosity fraction below 6%, we found three tidally elongated UDGs located close to massive galaxies. According to Wang2023, at the early steps of tidal disruption, nucleated dEs fall into the regime of UDGs and show tidal features such as an extended stellar envelope or tidal tails. Based on this statement, we add these three UDGs to the list of candidates, and show them in Fig. 9.